OpenStudy (anonymous):

What is the 7th term of the geometric sequence where a1 = -3 and a5 = -768?

OpenStudy (anonymous):

Select one: A. -12,288 B. -3,072 C. 12,288 D. 3,072

OpenStudy (anonymous):

\[a_1,a_2,a_3,a_4...\] \[a_1,a_1r,a_1r^2,a_1r^3...\] \[a_5=a_1r^4\] and so \[-768=-3\times r^4\]

OpenStudy (anonymous):

okie use of the formula ar^n-1 you have a5 as -768 so -768=-3(r)^5-1 divide 768 by 3 and then take 4th root put the value of r in the equation and find the final answer

OpenStudy (anonymous):

r = the common difference right?

OpenStudy (anonymous):

that makes \[r^4=\frac{-765}{-3}=256\]

OpenStudy (anonymous):

what @Danyboii said, take the 4th root to get \(r\) then compute \[-3\times r^6\]

OpenStudy (anonymous):

alright, one moment please.

OpenStudy (anonymous):

\(r\) is not the "common difference" that is for an "arithmetic sequence" this is a "geometric sequence" so \(r\) is the "common ratio"

OpenStudy (anonymous):

oh um... guys... you've lost me... how would I find r exactly? Or in other words, how would I find the common ratio? I'll give it a quick google and see if I can't teach myself.

OpenStudy (anonymous):

"to find the common ratio, divide the second term by the first term."

OpenStudy (anonymous):

yea but not in this case u r not given the second term

OpenStudy (anonymous):

Alright, so for this equation the second term is oh....

OpenStudy (anonymous):

that is why in this case you have to take the fourth root

OpenStudy (anonymous):

Oh ok I'm getting there (thank you both for being patient) So how on earth do I find this "fourth root"?

OpenStudy (anonymous):

just use the formula ar^n-1 so for a5 u have n as 5 a is -3 and = to -768 :3

OpenStudy (anonymous):

guess it

OpenStudy (anonymous):

or use a calculator

OpenStudy (anonymous):

lol 4\[\sqrt[4]{256}\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=4th+root+256

OpenStudy (anonymous):

ohhhhh god! Thank you guys... (math sucks for me). 12,288

OpenStudy (anonymous):

lol :p shouldnt it be in minus?

OpenStudy (anonymous):

should be negative

OpenStudy (anonymous):

yea since u r multiplying by -3

OpenStudy (anonymous):

oh lol oops I seriously got excited.

OpenStudy (anonymous):

you were there, you just forgot the \(-\)sign

OpenStudy (anonymous):

Thank you two so much you guys are great and smart and you two should be proud of yourselves for being so helpful and patient! You guys are the Gandhi's of Openstudy.

OpenStudy (anonymous):

Lol :p thanks :3 and yea ur welcome ;3

OpenStudy (anonymous):

yw